# SP-VIO: Robust and Efficient Filter-Based Visual Inertial Odometry with State Transformation Model and Pose-Only Visual Description

### v1

@@ Abstract
Due to the advantages of high computational efficiency and small memory requirements, filter-based visual inertial odometry (VIO) has a good application prospect in miniaturized and payload-constrained embedded systems. However, the filter-based method has the problem of insufficient accuracy. To this end, we propose the State transformation and Pose-only VIO (SP-VIO) by rebuilding the state and measurement models, and considering further visual deprived conditions. In detail, we first proposed a system model based on the double state transformation extended Kalman filter (DST-EKF), which has been proven to have better observability and consistency than the models based on extended Kalman filter (EKF) and state transformation extended Kalman filter (ST-EKF). Secondly, to reduce the influence of linearization error caused by inaccurate 3D reconstruction, we adopt the Pose-only (PO) theory to decouple the measurement model from 3D features. Moreover, to deal with visual deprived conditions, we propose a double state transformation Rauch-Tung-Striebel (DST-RTS) backtracking method to optimize motion trajectories during visual interruption.
Experiments on public (EuRoC, Tum-VI, KITTI) and personal datasets show that SP-VIO has better accuracy and efficiency than state-of-the-art (SOTA) VIO algorithms, and has better robustness under visual deprived conditions.
@@ I Introduction
Visual inertial odometry (VIO) technology is widely used in various mobile robots for autonomous navigation. Although different VIO algorithms have been developed in the past decade, roboticists still expect algorithms with lower computation costs, higher accuracy, and better robustness.
The mainstream VIO algorithms can be broadly divided into two categories: optimization-based [1, 2, 3, 4, 5] and filter-based methods [6, 7, 8, 9, 10]. Benchmark experiments [11, 12] show that filter-based VIO has the advantages of high computational efficiency and small memory requirements, and has a good application prospect in payload-constrained embedded systems. However, compared with the optimization-based VIO, filter-based methods are more efficient and also have the problem of insufficient accuracy, which limits the application of this method in some scenarios.
Taking MSCKF [10], the representative work of filter-based VIO, as an example, this method uses null space projection to simplify the residual model and then performs EKF update to estimate the corresponding pose. Because the 3D features used in the visual residual model are obtained only through a 3D reconstruction process and are not jointly optimized with the system’s pose, it is easy to cause the accumulation of linearization errors and lead to the decline of navigation accuracy. In addition, EKF is prone to filtering inconsistency problems, that is, misestimating the observability of the system leads to performance degradation, which is also an important factor affecting the accuracy of filter-based VIO [13].
@@
As mentioned above, an important problem of MSCKF-based VIO is that it is not completely decoupled from 3D features, which are still required to construct visual residuals. The newly proposed PO theory indicates that reprojection error can be obtained solely through camera pose, which is equivalent to traditional multi-view geometry description [15, 16]. Inspired by PO theory, we derive a visual residual model based only on pixel coordinates and relative pose, which is completely decoupled from 3D features to avoid being affected by inaccurate 3D reconstruction processes. Meanwhile, the new residual model does not need the simplification of left null space, and can be used for filter update directly.
@@
The above works are concerned with solving the accumulation of linearization errors caused by inaccurate system modeling. In addition, visual measurement information is also an important source of linearization errors. Due to the sensitivity of camera to the environment, it is easy to be disturbed in the actual motion, resulting in the visual tracking failure, which makes the system input discontinuous measurement information and affects the subsequent state estimation accuracy. The current common solution is the visual relocalization [29], but this method relies on matching with the existing localization scenes and does not work when the motion trajectory is not repeated, which limits the robustness of VIO under non-closed-loop motion trajectories. Rauch-Tung-Striebel (RTS) [30] is an efficient optimal backtracking smoothing method, which can optimize the cumulative error based on the historical state information, and is not limited by whether a closed-loop trajectory is formed. Therefore, we propose the DST-RTS backtracking smoothing method combining DST-EKF and RTS smoother to correct motion trajectories during visual interruption. This will improve the robustness of VIO system under visual deprived conditions.
@@
  • We propose a pose-only measurement model that can be used for filter-based VIO systems. This model is decoupled from 3D features and avoids the damage to accuracy caused by inaccurate 3D reconstruction while maintaining efficiency.
  • We propose a system model based on the double state transformation Kalman filter(DST-EKF), which enables the VIO system to obtain the improvement of robustness and consistency by defining the velocity error state and the position error state more strictly.
@@
  • A novel VIO algorithm named SP-VIO is proposed, which takes MSCKF [10] as the baseline system and integrates the above functions. We conduct comprehensive performance evaluations on both popular public datasets (EuRoC[31], Tum VI[32], Kitti Odometry[33]) and personal datasets. Experimental results show that SP-VIO has both high precision and robustness, with comprehensive performance better than SOTA VIO algorithms.
@@ II-C Filtering Inconsistency Problem
The current strategies to solve the inconsistency problem of filter-based VIO can be broadly classified into two categories. The first class of methods is to maintain the system’s observability consistent with the reality. Geneva [55] and Tong [56] combine the Schmidt Kalman filter (SKF) with the visual-inertial navigation system (VINS), which sets the Kalman filter gain corresponding to the unobservable state to zero, that is, the unobservable state is not corrected to ensure the consistency of the system. Hesch [20] developed an observability-constrained EKF (OC-EKF) for VIO that directly enforces the unobservable directions of the system in the update step of the EKF and showed that this significantly reduces filter inconsistency. Huang proposed the first-estimates Jacobian extended Kalman filter (FEJ-EKF)[17], which reduces the damage of the observability problem to the consistency of EKF. Li and Mourikis [13] applied the FEJ concept to an MSCKF design and showed that this improved consistency over a standard MSCKF. To solve the problem that FEJ is sensitive to poor initial estimates, yang proposed the FEJ2 method which is robust to large initialization errors[18].
@@ II-D Correction Strategy under Visual Deprived Conditions
III System Model Based on DST-EKF
ST-EKF has achieved good application results in the field of integrated navigation [21, 27, 28], but the observability analysis shows that only the state transformation of the velocity error state is not rigorous enough (see Subsection III-C). Therefore, we propose the DST-EKF method based on ST-EKF, and the state transformation of the position error state is also performed, aiming to further improve the accuracy of the system model definition.
III-A State Definition
@@
where ${}_{{b}_{k}}^{G}\boldsymbol{\phi}$ represents the error state of ${}_{{b}_{k}}^{G}\boldsymbol{q}$. Excepy for quaternion, other states can be used with the error definition (e.g. $\delta\boldsymbol{x}=\boldsymbol{\widetilde{x}}-\boldsymbol{x}$). The hamiton quaternion is used in this paper, and its extended additive error can be defined as:
@@
III-B Propagation and Augmentation
The traditional velocity error is defined as $\delta^{G}\boldsymbol{v}_{{b}_{k}}={{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}-{{}^{G}\boldsymbol{v}_{{b}_{k}}}$, only the difference numerical values is considered, and the difference between the coordinate systems is not considered. The velocity error redefined by the state transformation method is expressed as follows[21]:
$\delta_{ST}^{G}\boldsymbol{v}_{{b}_{k}}$ $={{}_{{b}_{k}}^{G}\boldsymbol{R}}{{}_{G}^{{b}_{k}}\boldsymbol{\widetilde{R}}}{{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}-{{}^{G}\boldsymbol{v}_{{b}_{k}}}$ (10)
$\approx\delta^{G}\boldsymbol{v}_{{b}_{k}}-[{{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}$
@@ III-B Propagation and Augmentation
$\delta_{ST}^{G}\boldsymbol{p}_{{b}_{k}}$ $={{}_{{b}_{k}}^{G}\boldsymbol{R}}{{}_{G}^{{b}_{k}}\boldsymbol{\widetilde{R}}}{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}-{{}^{G}\boldsymbol{p}_{{b}_{k}}}$ (11)
$\approx\delta^{G}\boldsymbol{p}_{{b}_{k}}-[{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}$
@@
$\delta_{ST}^{G}\boldsymbol{\dot{p}}_{{b}_{k}}$ $=\delta^{G}\boldsymbol{\dot{p}}_{{b}_{k}}-[{{}^{G}\boldsymbol{\dot{\widetilde{p}}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}-[{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\dot{\phi}}$ (12)
$\approx[^{G}\boldsymbol{p}_{{b}_{k}}\times][\boldsymbol{\omega}^{G}\times]{{}_{{b}_{k}}^{G}\boldsymbol{\phi}}+{\delta^{G}\boldsymbol{v}_{{b}_{k}}}$
$~{}~{}+[^{G}\boldsymbol{p}_{{b}_{k}}\times]{{}_{{b}_{k}}^{G}\boldsymbol{R}}{\delta\boldsymbol{b}_{{g}_{k}}}+[^{G}\boldsymbol{p}_{{b}_{k}}\times]{{}_{{b}_{k}}^{G}\boldsymbol{R}}{w_{g}}$
@@
$\boldsymbol{F}_{b}=\begin{bmatrix}-[\boldsymbol{\omega}^{G}\times]&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&-_{b}^{G}\boldsymbol{R}&\boldsymbol{0}_{3\times 3}\\ -[\boldsymbol{g}^{G}\times]\\ -[^{G}\boldsymbol{v}_{b}\times][\boldsymbol{\omega}^{G}\times]&-2[\boldsymbol{\omega}_{ie}^{w}\times]&\boldsymbol{0}_{3\times 3}&[^{G}\boldsymbol{v}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{R}_{b}^{w}\\ [^{G}\boldsymbol{p}_{b}\times][\boldsymbol{\omega}^{G}\times]&\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}&[^{G}\boldsymbol{p}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\end{bmatrix}$ (15)
$\boldsymbol{G}_{b}=\begin{bmatrix}-_{b}^{G}\boldsymbol{R}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ [^{G}\boldsymbol{v}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&{}_{b}^{G}\boldsymbol{R}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ [^{G}\boldsymbol{p}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{I}_{3\times 3}\end{bmatrix}$ (16)
@@
The state and covariance augmentation part is consistent with OpenVINS [6] and will not be repeated here. In addition, since we have modified the error state definition, the state update equation is modified as:
@@
III-C Observability Analysis
System observability plays a crucial role in state estimation [69, 70]. Understanding system observability provides a deep insight into the system’s geometrical properties and determines the minimal measurement modalities needed. The observability matrix can be construct as [20]:
$\boldsymbol{O}\triangleq\begin{bmatrix}\boldsymbol{H}_{0}{\boldsymbol{\Phi}_{0,0}}\\ \boldsymbol{H}_{1}{\boldsymbol{\Phi}_{1,0}}\\ \vdots\\ \boldsymbol{H}_{k}{\boldsymbol{\Phi}_{k,0}}\end{bmatrix}$ (18)
The nullspace ${\boldsymbol{N}}$ of EKF-based VIO should be of 4 d.o.f and relates to the global yaw and translation [20]:
${\boldsymbol{N}_{k}^{(EKF)}}=\begin{bmatrix}\boldsymbol{0}_{3\times 3}&{\boldsymbol{g}^{G}}\\ \boldsymbol{0}_{3\times 3}&{[{{}^{G}\boldsymbol{v}_{{{b}_{0}}}}\times]}{\boldsymbol{g}^{G}}\\ \boldsymbol{I}_{3\times 3}&{[{{}^{G}\boldsymbol{p}_{{{b}_{0}}}}\times]}{\boldsymbol{g}^{G}}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{I}_{3\times 3}&{[{{}^{G}\boldsymbol{p}_{{{f}_{l}}}}\times]}{\boldsymbol{g}^{G}}\end{bmatrix}$ (20)
As shown in (20), the global yaw observability of standard EKF is easily affected by the linearization error, so that ${\boldsymbol{N}}$ is reduced to only 3 d.o.f. This will cause the system to obtain the wrong observability and affect the filtering accuracy.
From [19], MSCKF and EKF-SLAM have equivalent observability, therefore, to facilitate observability analysis, we construct DST-EKF-SLAM. The new state vector ${\boldsymbol{x}_{k}^{(Ob)}}$ consists of the current inertial navigation state ${\boldsymbol{x}_{b}^{(Ob)}}$ and a set of feature position ${\boldsymbol{x}_{f}^{(Ob)}}$ :
${\boldsymbol{x}_{k}^{(Ob)}}$ $=\begin{bmatrix}{\boldsymbol{x}_{b}^{(Ob){T}}}&{\boldsymbol{x}_{f}^{(Ob){T}}}\end{bmatrix}^{T}$ (21)
${\boldsymbol{x}_{f}^{(Ob)}}$ $=\begin{bmatrix}{}^{G}\boldsymbol{p}_{{f}_{1}}^{T}&{\cdots}&{}^{G}\boldsymbol{p}_{{f}_{N}}^{T}\end{bmatrix}^{T}$ (22)
where the definition of ${\boldsymbol{x}_{b}^{(Ob)}}$ is consistent with (2), ${}^{G}\boldsymbol{p}_{{f}}$ represents the feature position expressed in the global frame $\{G\}$.
From Eq. (18) and (19), the right null space matrix of DST-EKF-SLAM can be obtained:
${\boldsymbol{N}_{k}^{(DST)}}=\begin{bmatrix}\boldsymbol{0}_{3\times 3}&{\boldsymbol{g}^{G}}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{I}_{3\times 3}&{[{{}^{G}\boldsymbol{p}_{{{b}_{k}}{{f}_{l}}}}\times]}{\boldsymbol{g}^{G}}\end{bmatrix}$ (23)
where ${{}^{G}\boldsymbol{p}_{{{b}_{k}}{{f}_{l}}}}={{}^{G}\boldsymbol{p}_{{{f}_{l}}}}-{{}^{G}\boldsymbol{p}_{{{b}_{k}}}}$
To show more clearly the improvement of DST-EKF consistency, we also derive the right null space of ST-EKF-SLAM:
${\boldsymbol{N}_{k}^{(ST)}}=\begin{bmatrix}\boldsymbol{0}_{3\times 3}&{\boldsymbol{g}^{G}}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{I}_{3\times 3}&{[{{}^{G}\boldsymbol{p}_{{{b}_{0}}}}\times]}{\boldsymbol{g}^{G}}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{I}_{3\times 3}&{[{{}^{G}\boldsymbol{p}_{{{b}_{k}}{{f}_{l}}}}\times]}{\boldsymbol{g}^{G}}\end{bmatrix}$ (24)
Compared with Eq. (23) to Eq. (24), DST-EKF-SLAM has the most concise right null space matrix about the global yaw, which is only related to the feature position and the gravity vector. Therefore, it is less affected by the linearization error in non-ideal cases, and has the most stable observability among the three algorithms.
IV Measurement Model Based on PO Theory

### v2

@@ Abstract
Due to the advantages of high computational efficiency and small memory requirements, filter-based visual inertial odometry (VIO) has a good application prospect in miniaturized and payload-constrained embedded systems. However, the filter-based method has the problem of insufficient accuracy. To this end, we propose the State transformation and Pose-only VIO (SP-VIO) by rebuilding the state and measurement models, and considering further visual deprived conditions. In detail, we first proposed the double state transformation extended Kalman filter (DST-EKF) to replace the standard extended Kalman filter (Std-EKF) for improving the system’s consistency, and then adopt pose-only (PO) visual description to avoid the linearization error caused by 3D feature estimation. The comprehensive observability analysis shows that SP-VIO has a more stable unobservable subspace, which can better avoid the inconsistency problem caused by spurious information. Moreover, we propose an enhanced double state transformation Rauch-Tung-Striebel (DST-RTS) backtracking method to optimize motion trajectories during visual interruption.
Monte-Carlo simulations and real-world experiments show that SP-VIO has better accuracy and efficiency than state-of-the-art (SOTA) VIO algorithms, and has better robustness under visual deprived conditions.
@@ I Introduction
Visual inertial odometry (VIO) technology is widely used in various unmanned systems for autonomous navigation. Although different VIO algorithms have been developed in the past decade, the unmanned systems industry still expect algorithms with lower computation costs, higher accuracy, and better robustness.
The mainstream VIO algorithms can be broadly divided into two categories: optimization-based [1, 2, 3, 4, 5] and filter-based methods [6, 7, 8, 9, 10, 11, 12]. Benchmark experiments [13, 14] show that filter-based VIO has the advantages of high computational efficiency and small memory requirements, and has a good application prospect in payload-constrained embedded systems. However, compared with the optimization-based VIO, filter-based methods are more efficient but also have the problem of insufficient accuracy, which limits their application in some scenarios.
Taking MSCKF [6, 7, 8, 9], the representative work of filter-based VIO, as an example, this method uses null space projection to simplify the residual model and then performs EKF update to estimate the corresponding pose. Because the 3D features used in the visual residual model are obtained only through a 3D reconstruction process and are not jointly optimized with the system’s pose, it is easy to cause the accumulation of linearization errors and lead to the decline of navigation accuracy. In addition, EKF is prone to filtering inconsistency problems, that is, misestimating the observability of the system leads to performance degradation, which is also an important factor affecting the accuracy of filter-based VIO [8].
@@
As mentioned above, an important problem of MSCKF-based VIO is that it is not entirely decoupled from 3D features, which are still required to construct visual residuals. The newly proposed PO theory indicates that reprojection error can be obtained solely through camera pose, equivalent to traditional multi-view geometry description [15, 16]. Inspired by PO theory, we derive a visual residual model based only on pixel coordinates and relative pose, which is completely decoupled from 3D features to avoid being affected by inaccurate 3D reconstruction processes. Meanwhile, the new residual model does not need the simplification of left null space, and can be used for filter update directly.
@@
The above works are concerned with solving the accumulation of linearization errors caused by inaccurate system modeling. In addition, visual measurement information is also an important source of linearization errors. Due to the camera’s sensitivity to the environment, it is easy to be disturbed in the actual motion, resulting in the visual tracking failure, which makes the system input discontinuous measurement information and affects the subsequent state estimation accuracy. The current common solution is the visual relocalization [29], but this method relies on matching with the existing localization scenes and does not work when the motion trajectory is not repeated, which limits the robustness of VIO under non-closed-loop motion trajectories. Rauch-Tung-Striebel (RTS) [30] is an efficient optimal backtracking smoothing method, which can optimize the cumulative error based on the historical state information, and is not limited by whether a closed-loop trajectory is formed. Therefore, we propose the DST-RTS backtracking smoothing method combining DST-EKF and RTS smoother to correct motion trajectories during visual interruption. This will improve the robustness of VIO system under visual deprived conditions.
@@
  • We propose a novel VIO system named SP-VIO. Specifically, we first propose the DST-EKF to replace the Std-EKF for improving the system’s consistency by more strict velocity and position error definitions. Meanwhile, we adopt PO visual description to decouple the measurement model from 3D features to avoid the linearization error caused by inaccurate feature triangulation.
  • We, for the first time, analyze the PO visual description based VIO system’s observability, and demonstrate that VIO system combining DST-EKF and PO-based measurement model have more stable unobservable subspace, thus better avoiding inconsistency problems due to linearization errors.
@@
  • We conduct comprehensive performance evaluations through Monte-Carlo simulations and real-world experiments, including popular public datasets (EuRoC[31], Tum VI[32], Kitti Odometry[33]) and personal datasets. Experimental results show that SP-VIO has both high precision and robustness, with comprehensive performance better than SOTA VIO algorithms.
@@ II-C Filtering Inconsistency Problem
The current strategies to solve the inconsistency problem of filter-based VIO can be broadly classified into two categories. The first class of methods is to maintain the system’s observability consistent with the reality. Geneva [55] and Tong [56] combine the Schmidt Kalman filter (SKF) with the visual-inertial navigation system (VINS), which sets the Kalman filter gain corresponding to the unobservable state to zero, that is, the unobservable state is not corrected to ensure the consistency of the system. Hesch [20] developed an observability-constrained EKF (OC-EKF) for VIO that directly enforces the unobservable directions of the system in the update step of the EKF and showed that this significantly reduces filter inconsistency. Huang proposed the first-estimates Jacobian extended Kalman filter (FEJ-EKF)[17], which reduces the damage of the observability problem to the consistency of EKF. Li and Mourikis [8] applied the FEJ concept to an MSCKF design and showed that this improved consistency over a standard MSCKF. To solve the problem that FEJ is sensitive to poor initial estimates, chen proposed the FEJ2 method which is robust to large initialization errors[18].
@@ II-D Correction Strategy under Visual Deprived Conditions
III Basic SP-VIO Framework
In this paper, the basic SP-VIO is developed based on OpenVINS[7], where the DST-EKF is used to replace the Std-EKF, and the measurement model based on Pose-only visual description is utilized to replace the original multi-view geometry based measurement model. The framework of SP-VIO algorithm is shown in Fig. 1.
III-A System Model Based on DST-EKF
III-A1 State Definition
@@
where ${}_{{b}_{k}}^{G}\boldsymbol{\phi}$ represents the error state of ${}_{{b}_{k}}^{G}\boldsymbol{q}$. Except for quaternion, other states can be used with the error definition (e.g. $\boldsymbol{x}=\boldsymbol{\widetilde{x}}+\delta\boldsymbol{x}$). The hamiton quaternion is used in this paper, and its extended additive error can be defined as:
@@
III-A2 Propagation and Augmentation
The traditional velocity error is defined as $\delta^{G}\boldsymbol{v}_{{b}_{k}}={{}^{G}\boldsymbol{v}_{{b}_{k}}}-{{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}$, only the difference numerical values is considered, and the difference between the coordinate systems is not considered. The velocity error redefined by the state transformation method is expressed as follows[21]:
$\delta_{ST}^{G}\boldsymbol{v}_{{b}_{k}}$ $={{}^{G}\boldsymbol{v}_{{b}_{k}}}-{{}_{{b}_{k}}^{G}\boldsymbol{R}}{{}_{G}^{{b}_{k}}\boldsymbol{\widetilde{R}}}{{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}$ (10)
$\approx-\delta^{G}\boldsymbol{v}_{{b}_{k}}+[{{}^{G}\boldsymbol{\widetilde{v}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}$
@@ III-A2 Propagation and Augmentation
$\delta_{ST}^{G}\boldsymbol{p}_{{b}_{k}}$ $={{}^{G}\boldsymbol{p}_{{b}_{k}}}-{{}_{{b}_{k}}^{G}\boldsymbol{R}}{{}_{G}^{{b}_{k}}\boldsymbol{\widetilde{R}}}{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}$ (11)
$\approx-\delta^{G}\boldsymbol{p}_{{b}_{k}}+[{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}$
@@
$\delta_{ST}^{G}\boldsymbol{\dot{p}}_{{b}_{k}}$ $=-\delta^{G}\boldsymbol{\dot{p}}_{{b}_{k}}+[{{}^{G}\boldsymbol{\dot{\widetilde{p}}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\phi}+[{{}^{G}\boldsymbol{\widetilde{p}}_{{b}_{k}}}\times]_{{b}_{k}}^{G}\boldsymbol{\dot{\phi}}$ (12)
$\approx-[^{G}\boldsymbol{p}_{{b}_{k}}\times][\boldsymbol{\omega}^{G}\times]{{}_{{b}_{k}}^{G}\boldsymbol{\phi}}+{\delta^{G}\boldsymbol{v}_{{b}_{k}}}$
$\quad-[^{G}\boldsymbol{p}_{{b}_{k}}\times]{{}_{{b}_{k}}^{G}\boldsymbol{R}}{\delta\boldsymbol{b}_{{g}_{k}}}-[^{G}\boldsymbol{p}_{{b}_{k}}\times]{{}_{{b}_{k}}^{G}\boldsymbol{R}}{w_{g}}$
@@
$\boldsymbol{F}_{b}=\begin{bmatrix}-[\boldsymbol{\omega}^{G}\times]&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&-_{b}^{G}\boldsymbol{R}&\boldsymbol{0}_{3\times 3}\\ [\boldsymbol{g}^{G}\times]\\ +[^{G}\boldsymbol{v}_{b}\times][\boldsymbol{\omega}^{G}\times]&-2[\boldsymbol{\omega}^{G}\times]&\boldsymbol{0}_{3\times 3}&-[^{G}\boldsymbol{v}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&-\boldsymbol{R}_{b}^{w}\\ -[^{G}\boldsymbol{p}_{b}\times][\boldsymbol{\omega}^{G}\times]&\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}&-[^{G}\boldsymbol{p}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\end{bmatrix}$ (15)
$\boldsymbol{G}_{b}=\begin{bmatrix}-_{b}^{G}\boldsymbol{R}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ -[^{G}\boldsymbol{v}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&-{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ -[^{G}\boldsymbol{p}_{b}\times]{{}_{b}^{G}\boldsymbol{R}}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 3}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{I}_{3\times 3}\end{bmatrix}$ (16)
@@
The state and covariance augmentation is consistent with OpenVINS [7] and will not be repeated here. In addition, since we have modified the error state definition, the state update equation is modified as:
@@
III-B Measurement Model Based on PO Visual description